OpenStudy (anonymous):

how to solve using elimination method... x-2y+4z=6 3y+9z=0 -2x+6y-z=-10

OpenStudy (john_es):

I would write the system into a matrix, \[\left[\begin{matrix} 1 & -2 & 4 & 6\\ 0 & 3 & 9 & 0 \\ -2 & 6 & -1 & -10\end{matrix}\right]\] Then use linear combinations of files.

OpenStudy (john_es):

For example, if you multiply the first row by 2 and sum first row with third row to substitute the third row, \[2F_1+F_3→F_3\] Then, \[\left[\begin{matrix} 1 & -2 & 4 & 6 \\ 0 & 3 & 9 & 0 \\ 0 & 2 & 7 & 2\end{matrix}\right]\]

OpenStudy (john_es):

And then you can do, \[2F_2-3F_3\rightarrow F_3\] \[\left[\begin{matrix} 1 & -2 & 4 & 6 \\ 0 & 3 & 9 & 0 \\ 0 & 0 & -3 & -6 \end{matrix}\right]\]

OpenStudy (john_es):

Now you only have to do backsubstitution in, \[ x -2y + 4z = 6 \\ 3y +9z= 0\\ -3z= -6 \]

OpenStudy (john_es):

\[z=2,y=-6,x=-14\]

OpenStudy (anonymous):

thank you so much for helping me to find the answer...

OpenStudy (john_es):

You're welcome.